# DiffX: Guide Your Layout to Cross-Modal Generative Modeling

### v1

@@ 4.4 DiffX Model in Modality-Shared Space
where $\mathcal{T}(\cdot)$ is a token selection operation that considers visual token positions of $\boldsymbol{T}$. The hyperparameter $\mu$ adjusts the weights of self-attention mechanism, whereas $\delta_{1}$ and $\delta_{2}$ are also two learnable scalars.
Compared with the Transformer blocks in SD, our DiffX-UNet incorporates Spatial-Transformer blocks that introduce an outer gated cross-attention layer and iterative gated self-attention layers. Moreover, we utilize the Long-CLIP caption embedding to enhance the layout features and the frozen cross-attention mechanism.
@@ 4.5 Training & Inference
Training. Algorithm 1 shows the optimization process.
1) The training of DP-VAE is critical prior to the training of DiffX for cross-modal generation, where we encode the cross-modal “RGB+X” image pair into the modality-shared latent space $\boldsymbol{z}$ and conduct end-to-end training. The training objective of DP-VAE includes the MSE loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016):
@@
2) The training objective of our DiffX model, namely the denoising function, can be formulated as:
@@
where $t\sim\text{Uniform}(\{1,...,T\})$ is sampled in each iteration. Then, the latent $\boldsymbol{z}_{t}\sim q(\boldsymbol{z}_{t}|\boldsymbol{z}_{0},\mathbf{M})$ is sampled from the forward process with aggregated task information, which is fed into the reverse diffusion model to predict the forward noise.
Inference. The sampling steps are identical to those of the conventional LDM, with the only difference being the dual-branch decoding by the dual-path decoders of DP-VAE. The main purpose is to estimate cross-modal data with features from the corresponding heads of the reverse model.
@@ 5 Experiments
5.1 RGB+X Datasets
We conduct experiments on three “RGB+X” datasets for four tasks to illustrate the generalizability of our DiffX. In detail, we adopt the clean-version FLIR dataset (F.A.Group 2019; Zhang et al. 2020) for Box $\rightarrow$ RGB+T and the MFNet dataset (Ha et al. 2017) for Segmentation (Seg.) maps $\rightarrow$ RGB+T. Meanwhile, the COME15K dataset (Zhang et al. 2021) is adopted for SOD $\rightarrow$ RGB+D and Edge $\rightarrow$ RGB+D.
@@ 5.3 Experimental Results and Comparison
Additionally, DiffX also enables the diverse cross-modal generation by editing the key captions, which is shown in Figure 6. In traffic-scene RGB+T generation, the generated weather condition can be changed by modifying keywords like “daytime” and “nighttime”, which is critical in autonomous driving systems. For the natural RGB+D generation, we can change the categories and attributes of the generated objects by adjusting the captions while keeping their structural integrity and outlines. It facilitates both creative manipulation and cross-modal image editing.
@@
Comparison with Existing Models.
As shown in Figure 7, we can see that our DiffX outperforms the baseline methods significantly. The VQVAE (Van Den Oord, Vinyals et al. 2017) exhibits poor performance due to its unsuitability for dual-path outputs in end-to-end training. The VQGAN (Esser, Rombach, and Ommer 2021) and DDPM (Ho, Jain, and Abbeel 2020) can generate overall good images, but lack precision in capturing target objects and details. The modified GLIGEN (Li et al. 2023), which also performs diffusion training in a modality-shared latent space, yields superior results compared with other baseline models. However, it fails to generate the accurate background in the nighttime scene despite using captions for guidance. It proves the effectiveness of employing Long-CLIP for caption embedding and incorporating cross-attention mechanism in our DiffX model.
On the other hand, we evaluate the generated image quality by LPIPS (Zhang et al. 2018) and FID (Heusel et al. 2017). Meanwhile, we adopt the pre-trained models in downstream tasks to evaluate the correspondence to the original layouts. We use EGFNet (Zhou et al. 2022) to test mean Intersection over Union (IoU) and mean Accuracy (Acc) on Seg. $\rightarrow$ RGB+T task, and use BBSNet (Fan et al. 2020) to test S-measure ($S_{\alpha}$) (Fan et al. 2017) and Mean Absolute Error ($M$) scores on Edge $\rightarrow$ RGB+D task. Additionally, we employ the SSIM score to evaluate the consistency of generated and ground-truth X images. The results in Table 1 demonstrate that our DiffX outperforms all baseline models in all metrics. Our impressive LPIPS and FID scores demonstrate the high-quality image generation by DiffX. Furthermore, the outstanding performance on downstream tasks highlights the effectiveness of DiffX in data augmentation for cross-modal vision tasks. Finally, the achieved SSIM scores on the X images (0.667 and 0.795 on the two tasks, respectively) demonstrate the advantage of our approach for consistent layout-aware generation.
@@
Firstly, we conduct the ablation study on the Laplacian Pyramid (LP) in our DP-VAE. Results in Table 2 show that the LP structure effectively improves the PSNR and SSIM (Wang et al. 2004), facilitating cross-modal image reconstruction through frequency-aware incorporation.
Secondly, we also convert DiffX into a standard generative model (denoted as DiffXs) to generate RGB and X separately. We aim to compare its unimodal generation with the cross-modal RGB+X generation by DiffX. Quantitative comparison in Table 3 shows that the cross-modal results generated by DiffX exhibit better LPIPS and FID scores. This is because the diffusion training in modality-shared space can provide key complementary information for RGB and X image generation respectively.
@@
Thirdly, we conduct the ablation study on the impact of utilizing text captions on SOD $\rightarrow$ RGB+D and Seg. $\rightarrow$ RGB+T tasks, which heavily rely on text captions. Figure 8 shows that DiffX can effectively capture the crucial captions, while the model without caption embeddings generates misaligned images, even affecting the image quality. It also necessitates incorporating Long-Clip for embedding long text captions in cross-modal generation.
@@
Given DiffX’s strong ability to generate cross-modal images, we also wonder if we can apply this framework to achieve more robust, controllable, and versatile generation across multiple modalities? Therefore, we conduct an experiment on COME15K dataset for “SOD $\rightarrow$ RGB+D+Edge” task to illustrate its adaptability. For this task, we modify the DP-VAE to support three-modality encoding and decoding, with the diffusion training carried out in a shared latent space that encompasses all three modalities.
@@
The example results are shown in Figure 9. It is evident that the DiffX model can effectively accommodate the “SOD $\rightarrow$ RGB+D+Edge” task by generating multiple modalities. DiffX can accurately generate diverse generative results that align well with the SOD map and provided caption, producing corresponding RGB images based on the layout conditions and caption while also matching depth images and edge images effectively. Therefore, we are confident in the adaptability of DiffX for generating cohesive data across wide-range modalities. In the future, if any datasets with well-aligned “RGB+T+D” images occur, DiffX is a standout choice for layout-guided multi-modal generation.
@@ 6 Conclusion
In this study, we introduce the DiffX model for cross-modal “RGB+X”generation under various layout conditions. The key innovation of DiffX lies in its use of diffusion-based training within a shared latent space across modalities. Extensive experiments have demonstrated the robustness and versatility of our method, also showcasing its adaptability to multi-modal generation. Despite a slight increase in computational costs (less than 2%), our method maintains superior efficiency in terms of both time and storage compared to training separate models for individual modalities. Moving forward, our goal is to enhance DiffX to support multiple-to-multiple generative modeling, enabling the generation of “RGB+X+Y” or even more modalities. We aim to incorporate diverse layout conditions and facilitate simultaneous multi-modal generation based on user requirements.

### v2

@@ 4.5 DiffX-UNet
where $\mathcal{T}(\cdot)$ is a token selection operation that selects the visual token positions of $\boldsymbol{T}$. The hyperparameter $\mu$ adjusts the weights of the SA mechanism, whereas $\delta_{1}$ and $\delta_{2}$ are also two learnable scalars. For the trainable FF layers, we also add the scaling factors with the tanh activation.
Compared with the conventional Transformer blocks in SD, our Spatial-Transformer blocks integrate the iterative gated SA layers to adaptively embed the joint-modality feature $\boldsymbol{H}^{*}$. Moreover, the text condition $\boldsymbol{c}^{*}$ embedded by Long-CLIP further enhances the long caption awareness.
@@ 4.6 Training & Inference
Training.
Algorithm 1 shows the optimization process.
1) The training of DP-VAE is critical prior to the training of DiffX for cross-modal generation, where we encode the cross-modal “RGB+X” image pair into the modality-shared latent space $\boldsymbol{z}$ and conduct end-to-end reconstruction. The training objective of DP-VAE includes the MSE loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016):
@@
2) The training objective of our DiffX model, namely the latent denoising function, can be formulated as:
@@
where $t\sim\text{Uniform}(\{1,...,T\})$ is sampled in each iteration. Then, the latent $\boldsymbol{z}_{t}\sim q(\boldsymbol{z}_{t}|\boldsymbol{z}_{0},\mathbf{M})$ is initialized, which is fed into the DiffX model to predict the forward noise.
Inference.
The difference in the sampling steps between our DiffX and conventional LDM is the dual-branch decoding by our DP-VAE. The main purpose is to estimate cross-modal images from features of the corresponding latent distributions of the DiffX model.
@@ 5 Experiments
5.1 Cross-Modal Datasets
We conduct experiments on three “RGB+X” datasets for five cross-modal generative tasks. In detail, we adopt the clean-version FLIR dataset (F.A.Group 2019; Zhang et al. 2020) for Box $\rightarrow$ RGB+T and the MFNet dataset (Ha et al. 2017) for Segmentation (Seg.) maps $\rightarrow$ RGB+T. Meanwhile, the COME15K dataset (Zhang et al. 2021) is adopted for SOD $\rightarrow$ RGB+D, Edge $\rightarrow$ RGB+D, and SOD $\rightarrow$ RGB+D+Edge.
@@ 5.3 Experimental Results and Comparison
Additionally, DiffX also enables the diverse cross-modal generation by editing the key captions, as shown in Figure 6. In traffic-scene RGB+T generation, the generated weather condition can be changed by modifying keywords like “daytime” and “nighttime”, which is critical in autonomous driving systems. For the natural RGB+D generation, we can change the categories and attributes of the generated objects by adjusting the captions while keeping their structural integrity and outlines. Therefore, it facilitates both creative manipulation and cross-modal image editing.
@@
Comparison with Existing Models.
As shown in Figure 7, we can see that our DiffX significantly outperforms the baseline methods in qualitative results. The VQVAE (Van Den Oord, Vinyals et al. 2017) exhibits poor performance due to its unsuitability for dual-path outputs in end-to-end training. The VQGAN (Esser, Rombach, and Ommer 2021) and DDPM (Ho, Jain, and Abbeel 2020) can generate overall good images but lack precision in capturing target objects and details. The modified GLIGEN (Li et al. 2023), which also performs diffusion training in a modality-shared latent space, yields superior results compared with other baseline models. However, it fails to generate accurate backgrounds in the nighttime scene despite using captions for guidance. It proves the effectiveness of employing Long-CLIP for caption embedding and joint-modality embedder for modal connection.
@@
On the other hand, we evaluate the generated image quality by LPIPS (Zhang et al. 2018) and FID (Heusel et al. 2017). Meanwhile, we adopt the pre-trained models in downstream tasks to evaluate the correspondence to the original layouts. We use EGFNet (Zhou et al. 2022) to test mean Intersection over Union (IoU) and mean Accuracy (Acc) on Seg. $\rightarrow$ RGB+T task, and use BBSNet (Fan et al. 2020) to test S-measure ($S_{\alpha}$) (Fan et al. 2017) and Mean Absolute Error ($M$) scores on Edge $\rightarrow$ RGB+D task. Additionally, we employ the SSIM score to evaluate the consistency of generated and ground-truth X images. The results in Table 1 demonstrate that our DiffX outperforms all baseline models in all metrics. Our impressive LPIPS and FID scores demonstrate the high-quality image generation by DiffX. Furthermore, the outstanding performance on downstream tasks highlights the effectiveness of DiffX in data augmentation for cross-modal vision tasks. Finally, the achieved SSIM scores on the X images (0.667 and 0.795 on the two tasks, respectively) illustrate the advantage of our approach for consistent layout-aware cross-modal generation.
@@
Firstly, we conduct the ablation study on the Laplacian Pyramid (LP) in our DP-VAE. Results in Table 2 show that the LP structure effectively improves the PSNR and SSIM (Wang et al. 2004), facilitating cross-modal reconstruction through frequency information incorporation.
Secondly, we aim to compare the unimodal generation with the cross-modal RGB+X generation by DiffX. Therefore, we convert DiffX into a standard generative model (denoted as DiffXs) to generate RGB and X separately. Quantitative comparison in Table 3 shows that the cross-modal results generated by DiffX exhibit better LPIPS and FID scores. This is because the diffusion and denoising training in modality-shared space can provide key complementary information for RGB and X image generation, respectively.
@@
Thirdly, we conduct the ablation study on the impact of text captions on SOD $\rightarrow$ RGB+D and Seg. $\rightarrow$ RGB+T tasks. The qualitative comparison in Figure 8 shows that DiffX can effectively capture the crucial captions, while the variant model without caption embeddings generates broken or misaligned images, ultimately affecting the image quality. Therefore, it also necessitates incorporating Long-CLIP for embedding long text captions in cross-modal generation.
@@
Given DiffX’s strong ability to generate cross-modal images, we also wonder if we can apply this framework to achieve more robust, controllable, and versatile generation across multiple modalities? Therefore, we conduct an experiment on COME15K dataset for “SOD $\rightarrow$ RGB+D+Edge” task to illustrate its adaptability. For this task, we modify the DP-VAE to support three-modality encoding and decoding, with the denoising training carried out in a shared latent space that encompasses three modalities.
@@
The example qualitative results are shown in Figure 9. It is evident that the DiffX model can effectively adapt to the “SOD $\rightarrow$ RGB+D+Edge” task. DiffX can accurately generate diverse modal images that align well with the provided SOD maps and captions, producing corresponding RGB images based on the layout conditions while also matching depth images and edge images effectively. Therefore, DiffX is a standout choice for generating cohesive data across wide-range modalities. We also believe it can extend to cross-modal video and 3D image generative modeling in modality-shared latent diffusion.
@@ 6 Conclusion
In this study, we introduce the DiffX model for cross-modal “RGB+X”generation under various layout conditions. The key innovation of DiffX lies in its use of diffusion-based training within a shared latent space across modalities. Extensive experiments have demonstrated the robustness and versatility of our method, also showcasing its adaptability to multi-modal generation. Despite a slight increase in computational costs (less than 1%), our method maintains superior efficiency and storage when compared to training separate models for generating individual modalities. Moving forward, our goal is to enhance DiffX to support multiple-to-multiple generative modeling, enabling the generation of “RGB+X+Y” or even more modalities. We also aim to incorporate diverse layout conditions and facilitate simultaneous multi-modal generation based on user instruction.